OpenStudy (anonymous):

Chapter 5:Indices and Logarithms @perl

OpenStudy (anonymous):

\[(-\frac{ 3 }{ 8 })^{-\frac{ 2 }{ 3 }}\]

OpenStudy (anonymous):

Find the value of the following.

OpenStudy (perl):

because of the negative exponent you can flip the contents of the fraction

OpenStudy (perl):

(-3/8)^(-2/3) = (-8/3)^(2/3)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[-\frac{ 8^{\frac{ 2 }{ 3 }} }{ 3^{\frac{ 2 }{ 3 }} }\]

OpenStudy (perl):

yes

OpenStudy (anonymous):

\[-\frac{ (2^3)^{\frac{ 2 }{ 3 }} }{ \sqrt[3]{3^2} }\]

OpenStudy (anonymous):

\[-\frac{ 4 }{ \sqrt[3]{9} }\]

OpenStudy (perl):

now you can rationalize denominator

OpenStudy (anonymous):

How to rationalize denominator?

OpenStudy (perl):

you have cube root of 3*3. if you have another cube root of 3 then you can rationalize denominator

OpenStudy (mathmath333):

a number is rationalized when u remove the square root, cube root,.... of it

OpenStudy (anonymous):

\[-\frac{ 4 }{ 3\times3 }\] it's like this?

OpenStudy (mathmath333):

no

OpenStudy (mathmath333):

\(\Large -\frac{ 4 }{ \sqrt[3]{9} }=-\frac{ 4 }{ \sqrt[3]{9} }\times \frac{\sqrt[3]{9}}{ \sqrt[3]{9} }\\ \Large =-\frac{ 4 \sqrt[3]{9}}{9}\)

OpenStudy (anonymous):

\[=\frac{ 4 }{ 9 }\]

OpenStudy (mathmath333):

i m feeling i made mistake

ganeshie8 (ganeshie8):

assuming we're interested only in real roots : \[\begin{align}\left(-\frac{ 3 }{ 8 }\right)^{-\frac{ 2 }{ 3 }} &= \left(-\frac{ 8 }{ 3 }\right)^{\frac{ 2 }{ 3 }}\\~\\ &= {\left[\left(-\frac{ 8 }{ 3 }\right)^2\right]}^{\frac{ 1 }{ 3 }}\\~\\ &= {\left[\frac{8^2}{3^2}\right]}^{\frac{ 1 }{ 3 }}\\~\\ &= {\left[\frac{(2^3)^2}{3^2}\right]}^{\frac{ 1 }{ 3 }}\\~\\ &= {\left[\frac{2^6}{3^2}\right]}^{\frac{ 1 }{ 3 }}\\~\\ &= {\left[\frac{2^6\times 3}{3^2\times 3}\right]}^{\frac{ 1 }{ 3 }}\\~\\ &= {\left[\frac{2^6\times 3}{3^3}\right]}^{\frac{ 1 }{ 3 }}\\~\\ &= \frac{(2^6\times 3)^{\frac{1}{3}}}{(3^3)^{\frac{1}{3}}}\\~\\ &=\frac{(2^6)^{\frac{1}{3}}3^{\frac{1}{3}}}{3}\\~\\ &=\frac{2^2\sqrt[3]{3}}{3}\\~\\ &= \frac{4\sqrt[3]{3}}{3} \end{align}\]

OpenStudy (anonymous):

The answer in the book say it's 4/9.

OpenStudy (ikram002p):

hmm

OpenStudy (ikram002p):

i agree o ganesh

OpenStudy (ikram002p):

i cant see any trick over there :O

OpenStudy (ikram002p):

like is it only dealing with reals or what :O

OpenStudy (anonymous):

Thnx @ganeshie8 @perl @mathmath333 @ikram002p